### v2

# 100,000 Crab giant pulses at 215 MHz detected with an SKA-Low prototype station

Abstract
We report detection and analysis of the largest ever low-frequency sample of Crab giant pulses (GPs) detected in frequency band 200 – 231.25 MHz. In total about 95000 GPs were detected, which, to our knowledge is the largest low-frequency sample of Crab GPs presented in the literature. The observations were performed between 2024-12-14 and 2025-03-31 with the Engineering Development Array 2, a prototype station of the low-frequency Square Kilometre Array telescope. The fluence distribution of GPs in the entire sample is very well characterised with a single power law N(F) Fα, where for all GPs, and and for GPs at the phases of the main pulse and low-frequency interpulse respectively. We do not observe flattening of the fluence distribution at the higher fluences. Although, the index of the power law fluence distribution remained approximately constant over the observing period, the normalisation of the distribution was strongly anti-correlated (coefficient ) with the scatter broadening time. The timescale ( weeks) of these variations indicates that intrinsic GP emission was modulated by the refractive scintillation as the signals propagated through the Crab Nebula and ISM. As a result, the measured fluence distribution was augmented for lower ( 2 ms) and diminished for higher ( 5 ms) scatter broadening time causing the GP detection rate to vary between 3000 and 100 per hour respectively (the correlation coefficient -0.9). Furthermore, for the first time at low-frequencies we observe indications of positive correlation (correlation coefficient 0.7) between the scatter broadening time () and dispersion measure. Our modelling favours the screen size pc with mean electron density e-cm-3 located within 100 pc from the pulsar (Crab Nebula or Perseus arm of the Milky Way galaxy). The observed frequency scaling of the scattering broadening time (where ) is in agreement with the previous measurements. The observed maximum spectral luminosities erg/Hz/s approach those of the weakest pulses from some repeating Fast Radio Bursts (FRBs). However, the distribution of pulse arrival times is consistent with a purely random Poisson process, and we do not find evidence of clustering. Overall, our results agree with the current views that GPs from extra-galactic Crab-like pulsars can be responsible for some very weak repeating FRBs, but cannot explain the entire FRB population. Finally, these results demonstrate an enormous transient science potential of individual SKA-Low stations, which can be unlocked by milli-second all-sky imaging.
keywords:
[…]
1 Introduction
Pulsar B0531+21, aka Crab, is one of the first discovered pulsars 1968Sci…162.1481S (85). It is also one of the most interesting and longest studied pulsars capturing the imagination of the general public due to the magnificent Crab Nebula. It is a truly fascinating object and one of the “living proofs” that supernova (SN) explosions can lead to formation of a neutron star as the Crab pulsar and the surrounding nebula were formed in the core-collapse SN 1054 observed by medieval astronomers. As such, the Crab pulsar is considered to be a young pulsar with a period ms. It is located at the distance of approximately 1.9 kpc 2023ApJ…952..161L (51) and moving with a velocity of about 120 km s-1 2008ApJ…677.1201K (38, 51).
The Crab pulsar is one of a very few pulsars emitting the so-called giant pulses (GPs). These are very short (s) bursts often exhibiting nanosecond substructure consisting of extremely narrow nano-shots (e.g., see 2007ApJ…670..693H (32) or 2010A&A…524A..60J (37)), which implies brightness temperatures reaching K 2003Natur.422..141H (33, 15). Crab GPs were discovered shortly after the discovery of the pulsar 1970Natur.226..529H (34, 84). Their energies follow a power law distribution 1972ApJ…175L..89A (3, 15, 10, 67, 63, 5) in contrast to the exponential or log-normal distribution of normal pulses observed in other pulsars 2012MNRAS.423.1351B (13), which indicates different emission mechanisms.
Interestingly, Crab average pulse profiles are considered to be an average of a large ensemble of GPs (e.g. 2021ApJ…920…38B (12, 60)). Hence, even though the physical mechanisms behind the GP emission remain unexplained (see review by 2016JPlPh..82c6302E (24)), they make the Crab pulsar a unique laboratory to study pulsar emission mechanisms (for example, see discussions in 2007ApJ…670..693H (32, 37, 12)).
Crab’s average pulse profile changes significantly as a function of frequency, and 8 components have been reported in literature (e.g. see 2023ApJ…959..111L (50) for a detailed summary). At frequencies below 2 GHz and at higher photon energies (from near infrared up to hard -rays) the average pulse profile consists of the main pulse (MP) with a precursor and a low-frequency interpulse (IP) at a phase offset of about 145° (about 13 ms) from the MP. At these frequencies the average pulse profiles are consistent with the standard model that MPs and IPs originate from physically very similar regions on the opposite sides of the neutron star (NS). However, this picture becomes much more complicated at frequencies 2 – 10 GHz. Firstly, at around 2.7 GHz the IP component disappears, while at higher frequencies (4 GHz) the IP re-appears, but shifted in phase by about 20° from the low-frequency IP. Therefore, it is believed to be a different component, and referred to as high-frequency IP. Furthermore, two additional high-frequency components (HFC1 and HFC2) show up at pulse phases following the IP (see Figure 1 in 2007ApJ…670..693H (32)). Finally, 2007ApJ…670..693H (32) discovered that properties (temporal and spectral structures, DM and polarisation) of high-frequency GPs occurring at phases of MP and high-frequency IP differ significantly. This implies that the physical mechanisms and conditions in the emission region of the high-frequency IP are very different to those of the MP - in contrast to standard pulsar model where their are expected to be nearly the same just on the opposite sides of the NS. Similarly, the high-frequency IP and additional components (HFC1 and HFC2) are most likely unrelated to the low-frequency IP and originate from different regions and physical processes. All these relatively recent findings demonstrate that, despite being one of the longest studied pulsars, Crab pulsar has still many mysteries, which can teach us a lot about the entire pulsar population. We note, however, that the micro and nano-second structure of Crab GPs cannot be observed at the low frequencies of the presented study due to the temporal scatter broadening caused by the ISM and Crab Nebula.
After the discovery of fast radio bursts (2007Sci…318..777L, 52, FRBs;), GPs from extra-galactic pulsars were hypothesised to be responsible for at least some types (i.e. repeating) of FRBs 2016MNRAS.457..232C (17, 58), and Crab GPs have been extensively studied in the context of FRBs (e.g. 2022NatAs…6..393N (66, 35)) as well as for testing this hypothesis (e.g. 2024ApJ…973…87D (21, 5, 63)). Based on the energy distribution of Crab GPs and other characteristics (e.g. time between pulses), these studies have disfavoured the theory that all FRBs are caused by GPs from extra-galactic pulsars, and it is more likely that majority of FRBs are caused by objects like magnetars 2019MNRAS.485.4091M (62, 27). However, some weaker FRBs, such as FRB 20200120E, could be caused by extra-galactic Crab-like pulsars (see 2022NatAs…6..393N (66)). For the most recent review of the FRB field see 2022A&ARv..30….2P (71), while FRB models are collated at https://frbtheorycat.org/.
While the discovery and early Crab GP studies (e.g. 1995ApJ…453..433L (54)) were performed at frequencies below 1 GHz, the majority of later investigations were conducted at frequencies above 1 GHz (e.g. 2024ApJ…973…87D (21, 5, 22, 10)) up to even 15.1 GHz (e.g. 2010A&A…524A..60J (37, 32)). On the high end of electromagnetic spectrum GP studies extend to -rays; see 2021Univ….7..448A (2) for the most recent review, and 2021Sci…373..425L (49, 1) for recent results at PeV energies.
The advent of new (LOFAR 2013A&A…556A…2V (87), MWA 2013PASA…30….7T (86, 90), LWA 2013ITAP…61.2540E (25) etc.) and upcoming instruments, like low-frequency Square Kilometre Array telescope (2009IEEEP..97.1482D, 19, SKA-Low;)111https://www.skatelescope.org/ with large fields-of-view (FoVs), has begun to unlock the wealth of information available at these frequencies, and reinvigorated pulsar studies at frequencies below 400 MHz. Several recent GP studies were performed down to 110 MHz (e.g. 2012A&A…538A…7K (39, 63, 68, 64, 11)). They addressed various aspects of Crab GPs ranging from monitoring plasma in the surrounding environment and along the line of sight (LoS) to the pulsar, through variations in dispersion measure (DM) and scatter broadening time to statistical properties of GPs, their emission mechanisms and similarities to FRBs.
This paper presents results from one such instrument, the Engineering Development Array version 2 (2022JATIS…8a1010W, 89, EDA2;) a prototype station of the SKA-Low. Despite modest sensitivity we were able to record the largest ever sample (95000) of Crab giant pulses observed at SKA-Low frequencies, superseding even the largest ever sample of 30000 Crab GPs recorded by 1995ApJ…453..433L (54). The EDA2 was originally deployed as a risk mitigation and reference system to the official SKA-Low prototype station the Aperture Array Verification System 2 station (2022JATIS…8a1014M, 59, AAVS2;). Real-time station beamforming and pulsar observing mode were commissioned at an early stage and already resulted in early SKA-Low pulsar science results 2022PASA…39…42L (48, 47). Both stations have also been used for all-sky transient searches 2021PASA…38…23S (80).
Recently, a new pipeline, based on a code developed for the MWA by 2024PASA…41…11S (78), has been implemented and deployed for non-targeted (“blind”) searches and monitoring of repeating FRBs, magnetars, RRATs and pulsars with a single station beam of EDA2. In order to test and verify the pipeline, multiple observations of Crab pulsar were performed, and led to detection of thousands of GPs. These detections have confirmed that the station and the pipeline are capable of detecting highly dispersed pulses, and can be successful in searching for and/or monitoring FRBs, magnetars, RRATs or other fast transients. Especially, once the full FoV of the stations is unlocked by multi-beaming or high-time resolution all-sky imaging as proposed by 2022aapr.confE…1S (83).
Moreover, the huge sample of Crab giant pulses turned out to be extremely valuable for scientific analysis. We have used this sample to perform statistical analysis of the Crab GPs, and as a demonstration of the potential of low-frequency telescopes for monitoring plasma in local environments and/or along the LoS towards pulsars, FRBs and other objects. Additionally, we have, once more, examined the possibility of extra-galactic Crab-like pulsars being FRB progenitors.
This paper is structured as follows. In Section 2 we describe the EDA2 telescope and summarise the Crab observations performed for this paper. Section 3 describes the pipeline and data processing, while Section 4 summarises data analysis. Results are presented in Section 5, and discussed in Section 6. Finally, in Section 7 we summarise, make conclusive remarks and discuss future directions and prospects.
2 EDA2 telescope and observations
The data for this study were recorded between 2024-12-14 and 2025-03-31 with the EDA2 station. The details of EDA2 can be found in the description paper 2022JATIS…8a1010W (89), and only a brief summary relevant to this project is provided here. The EDA2 has been a risk mitigation and a reference to the AAVS2 station, with the only major difference being the antenna design. The EDA2 comprises 256 MWA bow-tie dipoles, and was used as a reference to AAVS2 station (already de-commissioned) composed of 256 SKALA 4.1 antennas. Both instruments were located close to each other (some 500 m apart) at the SKA-Low site, Inyarrimanha Ilgari Bundara observatory222https://www.csiro.au/en/about/facilities-collections/mro. Analogue signals from individual antennas are digitised and the entire bandwidth (400 MHz) channelised into 512 coarse channels in the backend Tile Processing Units (2017JAI…..641014N, 65, TPMs;). One of the main tasks of TPMs is to form real-time station beams in 1.08 usec time resolution over the entire SKA-Low observing bandwidth (50 – 350 MHz).
For this project, a single station beam and 40 coarse channels covering the frequency band 200.00 – 231.25 MHz (i.e. observing bandwidth BW31.25 MHz) were recorded. Observations with wider bandwidths are also possible, and the system has been confirmed to be stable at least up to a BW100 MHz. During each observing session 1 h (all observations except 3) or 1.5 h (3 observations) of data were recorded and centred on transit time of the Crab pulsar. The recorded station beam complex voltages were saved to DADA format files with complex voltages from each coarse channel saved to a separate file (i.e. 40 DADA files per observation in total).
3 Data processing pipeline
The recorded DADA files have been processed with eda2frb pipeline333https://github.com/marcinsokolowski/eda2frb developed for this project (Fig. 1). The pipeline has been based on the FRB search pipeline applied to MWA incoherent beam 2024PASA…41…11S (78)444athttps://github.com/marcinsokolowski/mwafrb. It has been executed on a server computer installed at the observatory and consists of the following main steps. First, complex voltages in each coarse channel DADA file are fine channelised into 64 fine channels (approximately 14.5 kHz each) by a custom-developed spectrometer using FFTW (CPU) or cuFFT (GPU) libraries (the choice is optional). The resulting spectra (in 69.12 usec time resolution) are averaged to 0.96768 ms time resolution, and saved to separate filterbank files. Then, 40 filterbank files from each coarse channel are merged into a single filterbank covering the entire observing bandwidth (31.25 MHz). These steps are represented by blue boxes in Figure 1.
The single pulse search is performed on the resulting wideband filterbank file using both PRESTO 2011ascl.soft07017R (73) and FREDDA 2019ascl.soft06003B (4) software packages. One of the PRESTO executions was performed at Crab pulsar’s DM=56.72 pc/cm3 as measured during this period by the Jodrell Bank observatory555https://www.jb.man.ac.uk/pulsar/crab/crab2.txt. Selected FREDDA candidates were visually inspected, and an example of a GP with the highest signal-to-noise (SNR) ratio of 354 recorded in this campaign is shown in Figure 2. This is the brightest pulse in the sample with flux density of 16.6 kJy and fluence 76 Jy ms. We note that the pulse broadening effects were not removed in our analysis (e.g. via de-convolution, 2003ApJ…584..782B (7)), which may lead to a factor of 2 underestimation of the intrinsic peak flux density (see discussion in 2007ApJ…665..618B (11)).
It is worth mentioning that some previous studies recorded low-frequency (600 MHz) GPs upon receiving a trigger from high-frequency (1400 MHz) detections (e.g. 1999ApJ…517..460S (77)). However, in our study observations were performed at a single frequency (31.25 MHz bandwidth centred on 215 MHz), which was not strongly impacted by temporal pulse broadening, ionospheric scintillation or RFI 2016ApJ…829…62E (23). The moderate data rates ( 0.5 TB/h) enabled us to record 1 – 2 h observations. Hence, we were able to identify GPs independently (without relying on any external information) by performing a single pulse search offline using the standard software packages (i.e. PRESTO and FREDDA).
4 Data analysis
4.1 Post-processing of PRESTO output
The output from PRESTO single_pulse_search.py consists of three types of files: dat binary files contain de-dispersed time series as a series of float values, inf text files contain metadata, and singlepulse files contain a list of identified single pulse candidates (including arrival time, DM etc.). The following additional steps were performed on PRESTO output in preparation for the later analysis:
Merging candidates - a single GP may result in multiple candidates with different arrival times, DMs and SNRs. Therefore, in order to avoid double-counting, the candidates were merged into a single candidate with maximum SNR within a 45 ms time window as highly scattering-broadened pulses can extend up to this duration. In principle, we could have missed two GPs within one pulsar period (33 ms), but the results of merging were visually inspected and such cases have never been found. In fact, so far, the so-called double (or “twin”) GPs, from MP and IP within the same pulsar period, have only been reported in L-band 2010A&A…515A..36K (40). Also, recently, 2023ApJ…959..111L (50) reported detections of clusters of micro-bursts (even up to 6) within a single pulse window (1 ms), but such instances cannot be identified in our data due to scatter broadening at our frequencies and 1 ms time resolution.
In order to find double-GPs, we performed a dedicated search in every recorded de-dispersed time series, and did not find any instance of two pulses separated by less than a pulsar period (33 ms). Hence, we are confident that the merging procedure does not affect the statistical analysis of fluence and flux density distributions, or any other parts of the presented analysis.
Extraction of de-dispersed time series - The de-dispersed time series for each dataset were extracted from the PRESTO binary files (dat files) generated by the de-dedispersion at DM=56.72 pc/cm3. First, a running median and standard deviation (derived from the interquartile range) were calculated. The running median value was then subtracted as a baseline and the resulting time series was divided by the running standard deviation, which converted the uncalibrated flux units of the timeseries into SNR units. The resulting SNR-timeseries was flux-calibrated in Jy using the procedure described in Section 4.2, and flux density, fluence and SNR of each PRESTO candidate (after merging) were saved to the output text file. The timeseries extraction code is available at https://github.com/marcinsokolowski/presto_tools.
4.2 Flux Density Calibration
The de-dispersed time series converted to SNR units (Section 4.1) and SNRs reported in the .singlepulse log files were flux-calibrated using equation 1. If is the uncalibrated flux density measured in the data (de-dispersed timeseries), the calibrated flux density was calculated as:
where and are standard deviation of the noise as expected from the simulation and measured in data respectively, and ratio is the calibration factor (or inverse of the gain). The expected standard deviation of the noise () can be calculated from the radiometer equation:
where and are observing bandwidth and integration time respectively. is the System Equivalent Flux Density (SEFD) for Stokes I polarisation (hence and this factor is not included under the square root). For each dataset, was calculated using the SKA-Low-sensitivity calculator 2022PASA…39…15S (79)666https://github.com/marcinsokolowski/station_beam, which provides very accurate SEFD values as verified by extensive tests during the EDA2 and AAVS2 commissioning and verification stages 2022JATIS…8a1010W (89, 82). The mean value during the Crab pulsar transit at elevation 41°was Jy (for reference EDA2 SEFD at zenith is 2500 Jy). Hence, the expected noise was Jy for the observing parameters Hz (=31.25 MHz), sec (1 ms).
The flux-calibrated timeseries (in Jy) was used to calculate calibrated fluence (in units of Jy ms) as a sum (i.e. discrete integral) of the flux density under the pulse. For each GP reported by PRESTO (after merging), the fluence was calculated as the following sum:
where is the time bin width (1 ms), is the flux density in the bin , and the sum extends over the positive flux density bins surrounding the peak flux. The sum includes all positive flux density values before and after the flux density peak. Hence, the summation stops when the first negative value is encountered before and after the peak flux density bin, which indicates that the flux density reached the level consistent with noise. Negative values are present because of the subtraction of the running median as a baseline.
After calibration of each dataset (typically one hour observation per day), all parameters (arrival time, SNR, calibrated flux density, calibrated fluence etc.) describing every GP detected by PRESTO (after merging) were saved to a single text file (calibrated_pulses.txt). This file was a starting point for further statistical analysis. The scripts used for this processing are publicly available777https://github.com/marcinsokolowski/crab_frb_paper.
5 Results
The text file calibrated_pulses.txt served as a starting point for further analysis. The completeness threshold was verified to be SNR10, which at low scattering conditions (2 ms) corresponded to flux density threshold 500 Jy, and fluence threshold 1400 Jy ms. A scatter-plot (not included here) of all GP SNR vs. fluence confirmed that F1400 Jy ms above this value there are no GPs approaching the PRESTO threshold SNR5 which would not be detected (even during high scattering period when peak flux density may be reduced). Hence, no GPs with fluence F1400 Jy ms were missed (i.e. the sample is complete).
5.1 Fluence distribution
The GPs from the main pulse (MP) and interpulse (IP) were separated by calculating GP phase (using time of arrival and pulsar period) and matching the phase of MP and GP in the average profile. The fluence distributions of GPs associated with MP and IP were separately plotted and fitted with a power law (Figure 9):
where is the normalisation corresponding to number of counts at the reference value (1000 Jy ms for the fluence and 1000 Jy for the flux density distributions respectively). Therefore, the fitted value of can be interpreted as an hourly rate of GPs with these respective values of fluence or flux density.
Approximately 85% of the detected GPs occur at phase of the MP (the remaining 15% at phase with IP), which is in very good agreement with 86% measured at similar frequencies with MWA and 84% at 732 MHz 2017ApJ…851…20M (63). It is also very close to 87% at 1200 MHz 2012ApJ…760…64M (64).
The power law index fitted to separate datasets was approximately constant as a function of time (). On the other hand, the fitted power law normalisation () was strongly anti-correlated (coefficient 0.9 and p-value ) with the scatter-broadening time (). Hence, was higher when the was lower and vice versa (Fig. 7). This resulted in the fluence distribution “floating” up and down under the low (2 ms) and high (5 ms) scattering conditions respectively. Consequently, the number of detected GPs was also strongly correlated (coefficient 0.9) with the scatter broadening time (compare Figures 6 and 3) and increased and decreased when the scattering time was lower and higher respectively (Fig. 6).
Interestingly, our results confirm (probably yet another time) that PRESTO is a very good “fluence” detector as higher scattering times did not compromise the detection efficiency. This is based on the observation that the turn-over fluence , which is a fluence below which the number of detected GPs starts to decrease due to missing fainter GPs, remains approximately constant for various scatter broadening times. A weak correlation between and was indeed observed, indicating that as expected, higher scattering time reduces peak flux density resulting in fewer fainter pulses detected (i.e. is higher for higher values). However, this correlation is much weaker than could be expected for purely peak flux based detection software, while PRESTO implements boxcar convolution which makes it much closer to a perfect “fluence detector”.
5.2 Scatter broadening and DM variations
The scattering time was measured using GPs with SNR10. Each of these high-SNR pulses was fitted as a Gaussian pulse with an exponential tail (equation 15 in A), which yielded scatter broadening time (e.g. Figure 2). For each dataset all SNR10 GPs were fitted, and the mean and standard deviation of the resulting scattering time distribution provided the mean scattering time and its error for a particular dataset (i.e. date).
Scattering time and DM as a function of time are shown in Figure 3, where DM measured by Jodrell Bank 1993MNRAS.265.1003L (55)888http://www.jb.man.ac.uk/~pulsar/crab/crab2.txt is also plotted as a reference. A systematic constant offset between DM measured in this work and Jodrell Bank measurements is noticeable (see also Figure 17 and discussion in B). Especially, at MJDs between the start and 60680 days, before the increase in scattering time begun, which could affect our low-frequency timing-based DM measurements. Such DM offsets are not uncommon for measurements performed at different frequencies (this work at 215 MHz and Jodrell Bank in L-band), and may result from chromatic DM (see, for example2019ApJ…882..133K (42, 16)). However, this does not affect our analysis.
The DM was measured using standard pulsar timing performed with the TEMPO2 software package using frequency dependent pulse profiles. These measurements may be impacted (overestimate DM) by the changes of the average pulsar profile due to scattering - mainly by shifting the peak and changing the shape of the average pulse profile. Therefore, as a cross-check the DM was also measured by fitting profiles of individual GPs and finding DM maximising their peak flux density (i.e. SNR). The comparison of the DM measured with this method and timing is shown in Fig. 17 in B, where we also discuss other tested methods of measuring DM from individual GPs. The DM values measured from individual GPs are similar to the results of the timing analysis, and confirm that to the first order the latter are correct and the increase in DM is of physical origin (not an artefact of analysis). Although, during high scatter broadening period (between MJD 60700 and 60740), the increase in DM may be slightly lower than shown by the red filled circles in Figure 3, the DM values are lower bound by the well-established Jodrell Bank measurements (red stars). This ensures that our further analysis is valid to within a factor of 2 (in the worst case).
There are indications of positive correlation between DM and scattering time (Fig. 5), and as the scattering time gets higher DM also increases and vice versa (the Pearson correlation coefficient 0.7 and the p-value ). However, we note that this correlation may be slightly exaggerated by the earlier-mentioned impact of scattering on timing-based DM measurements. A similarly strong correlation of scatter broadening time and DM (correlation coefficient 0.560.01) based on normal pulses was observed in Crab pulsar at 610 MHz by 2018MNRAS.479.4216M (61), and at 350 MHz by 2019MNRAS.483.1224D (22). Hence, our measurements are consistent with the earlier observations of scatter broadening and DM correlations at higher frequencies. Nevertheless, we note that more data and robust DM measurements methods are required to confirm the correlation in the presence of strong scatter broadening.
The lag correlation of DM and scatter broadening time variations yielded a lag consistent with zero (to within a daily cadence of our observations). This is consistent with the results reported by 2018MNRAS.479.4216M (61), but significantly differs from the 30 days lag observed by 2008A&A…483…13K (45). It seems that the DM/ increase event observed by 2008A&A…483…13K (45) was caused by a plasma screen with a different structure and/or geometry of the Earth, pulsar, screen.
Finally, we have also compared scatter broadening times with earlier observations. Thus, the measured scatter broadening times were scaled to 610 MHz with a scaling factor as measured for the Crab pulsar 2019ApJ…874..179K (44, 11, 63). The resulting distribution is very similar and only slightly shifted towards lower scatter-broadening times with respect to the distribution in Figure 3 in 2018MNRAS.479.4216M (61). This indicates that the scattering conditions were likely similar during both events, and possibly caused by similar plasma structures.
5.3 Number of GPs vs. scatter broadening
One of the most interesting aspects originally noticed during the first observations in December 2024, were large variations in the numbers of detected GPs between 1-hour observations on different days (i.e. the rate of detectable GPs varied). During the first night of observations (2024-12-14) about 1400 GPs were detected by PRESTO, while on 2024-12-25 it reached 3000 to drop down to around 250 in the middle of February 2025 (Fig. 4). Interestingly, it started to increase again at the end of February 2025 to reach 2000 at the end of March (on 2025-03-27). The number of detected GPs was strongly correlated with the scatter broadening time (compare Figures 3 and 4, and see correlation in Figure 6).
The number of detected GPs as a function of scattering time has been modelled (black curve in Figure 6) using the scatter broadening time dependent fluence distribution (the fitted linear function in Figure 7), and functional form of pulse profile. Equation 15 in A was used to model giant pulse profiles as a function of time for different values of scatter broadening time with other parameters fixed (, , and ms). As discussed in Section 5.3 the fluence distribution of GPs depends on scatter broadening time. Hence, we used the linear dependence as fitted in Figure 7 to calculate normalisation of the fluence distribution for a specific . Then the expected number of GPs for particular value of was calculated as an integral of the fluence distribution above the minimal fluence . This minimal fluence was calculated as the fluence corresponding to the SNR threshold () used by the detection software PRESTO (SNR=5, i.e. peak flux density = 5230 Jy ms). The number of GPs as a function of and the results of our modelling are shown as a black curve in Figure 6. We note that this is only an approximation, because as discussed earlier by using convolution with boxcar function, PRESTO effectively implements a fluence detector, and this has not been reflected in our model. However, even this simple model reproduces the observed dependence very well. This gives us confidence that the variations in number of detected GPs can be explained by the up and down variations of the fluence distribution caused by refractive scintillation, which is also supported by the weeks timescale of these variations. Similar, variations in GP rates were observed by 1995ApJ…453..433L (54) in their largest ever recorded sample of GPs, which they also attributed to propagation effects. The explanation in terms of refractive scintillation is discussed in Section 6.1.2.
5.4 Flux density, SNR and spectral density distributions
Since, flux density was calculated from SNR using equation 1, these two quantities are related by a linear relation and are expected to follow the same distribution. The left image in Figure 8 shows the distribution of the peak flux density of GPs in the entire sample. The distribution is well described by a power law fitted above the completeness threshold of 500 Jy. The resulting power law index is , while the fit to the SNR distribution yielded a very similar value of .
Using the calibrated peak flux density and the distance to the pulsar D=2 kpc 2023ApJ…952..161L (51, 38), the peak spectral luminosity was calculated according to the following equation:
where is the peak flux density of the pulse. The resulting distribution is shown in the right panel of Figure 8. The maximum values are Lerg/s/Hz, which is consistent with the Crab spectral luminosities reported in literature (e.g. compare with Fig. 3 in 2022NatAs…6..393N (66) or Fig. 1 in 2018NatAs…2..865K (43)). These maximum spectral luminosities are a 2 – 3 orders of magnitude lower than the weakest repeating FRBs. However, assuming that there is enough energy in the system, the wait time for erg/s/Hz pulse is relatively short (of the order of 50,000 years). Hence, extra-galactic Crab-like pulsars may be responsible for weak repeating FRBs from the local Universe (see discussion in Section 6.2.1). However, the observed spectral luminosities are more than 10 orders of magnitude lower than typical or brightest FRBs. Hence, it is highly unlikely that such FRBs can be produced by objects similar to Crab pulsar (see discussion in Section 6.2) as there is not enough energy available. Additionally, based on our spectral luminosity distribution, the waiting time for such bright pulses (erg/s/Hz) is more than 10 orders of magnitude longer than the Hubble time.
6 Discussion
In this section, we discuss modelling of the observed variability of DM, and scatter broadening time as a function of time, and its physical interpretation. We also interpret our results in the context of FRBs and re-examine the possibility of extra-galactic Crab-like pulsars being FRB progenitors.
6.1 Modelling DM and variations
The DM and variations have been modelled as a plasma screen (or a “blob”) passing through the line of sight between the Earth and the Crab pulsar (Figure. 10).
We consider two distances between Earth and the screen (): (i) standard case midway (1000 pc) between the Earth and pulsar, and (ii) close to the pulsar (some 1900 pc from Earth) and consistent with the Crab Nebula or more broadly Perseus arm of the galaxy (explanation will be provided in Section 6.1.2). These two cases correspond to the ratio x= (Fig. 10) equal to 1 for (i) and 20 for (ii) respectively. For x=1 the velocity of the scintillation pattern with respect to observer on Earth is approximately equal to the velocity of the pulsar (eq. 10), while it can be higher for due to “lever-arm” argument (see equation 10 or discussion in 1995ApJ…451..717G (30)).
6.1.1 Modelling DM variations
The DM variations shown in Figure 3 have been modelled as a sphere of electron plasma passing transversely through the LoS between the pulsar and Earth (Fig. 10) using the following equation:
where the free parameters of the fit were additional small DM offset, the amplitude of the DM peak, the timescale of the transition of the plasma sphere through the LoS. These free parameters are independent, and they can be expressed as the following physical parameters: , and , where is the spatial standard deviation of the 2D Gaussian distribution describing the electron density in the plasma cloud, and is the cloud’s velocity (actual or apparent). The resulting parameters that minimised per degree of freedom (i.e. reduced ) were: pc cm-3, days, days, and pc cm-3.
Based on the fitted Gaussian parameters, physical parameters of the cloud can be estimated. The timescale of the entire DM-increase event is T 75 days, and the mean can be calculated as:
Assuming that the transverse motion of the screen relative to the observer on Earth is dominated by the pulsar’s proper motion =120 km/s (see eq. 10 and discussing in Section 6.1.2), we estimate the size of the cloud to be R pc, and the resulting mean electron density in the cloud to be 380 e-cm-3.
Based solely on our data, it is not possible to measure , , and without additional assumptions. Therefore, we have inferred physical parameters for the case of apparent transverse velocity of the screen km/s, and the results are summarised in the middle part of Table 1. We note that and for other values of can be calculated by scaling results for km/s. For example, and for km/s can be calculated by respectively dividing and multiplying by a factor 120/60.
The characteristic spatial scale of the plasma blob resulting from the fit is pc, and pc is consistent with the cloud size R pc estimated above using the mean values. This value leads to peak electron density 2600 e-cm-3. As an additional check, we performed a standard DM structure function analysis (for description see Section 5.2 in 2018MNRAS.479.4216M (61)) and obtained a characteristic timescale of 12.9 days, which corresponds to pc (when multiplied by 120 km/s) and agrees nearly perfectly with the value.
Hence, all tested methods lead to consistent results, and the resulting parameters are physically plausible. Despite the different epochs of observations, frequencies etc., they are also in-line with the studies by 2018MNRAS.479.4216M (61, 45, 29) as summarised in Table 1. This suggests that similar scintillation/scattering events due to plasma screens of the size pc and electron densities of the order of – are relatively common for the Crab pulsar. The observed electron densities are consistent with the filamentary structures inside the Crab Nebula (e.g. 1957PASP…69..227O (69, 26)).
[…]
| DM [pc cm-3] | 0.02 | 0.03 | 0.10 | 0.15 |
| Cloud size based on a | ||||
| [e- cm-3] | 380 | – | Not calculated | |
| Modelling DM(t) variations | ||||
| Peak [/cm-3] | 2500200 | |||
| Modelling (t) variations | ||||
| Range of [ms] | 2 – 5 | 8 – 27 | 0.1 – 0.7b | Not measured |
| Range of [ms] | 2 – 5 | 0.8 – 2.5 | 2.5 – 15.6b | Not measured |
| scaled to 215 MHz with | ||||
Table notes:
a
Details in Section 6.1.1.
b
The baseline level of scattering time in this work was around 0.1 ms which corresponds to 2.5 ms at 215 MHz. However, sudden drops down to 0 ms were also observed (see Figure 2 in 2018MNRAS.479.4216M (61)). Since we are not sure if these were valid measurements of extremely low scattering time we used the baseline level as a reference.
6.1.2 Effects of refractive scintillation
The long observed timescale 70 days of the variations in GP rate and fluence distribution is a manifestation of intensity modulations caused by refractive scintillation. Moreover, estimates of the size pc of the plasma cloud agree very well with the large-scale structures required by refractive scintillation (see, for example 1999ApJ…514..272B (9)). The refractive scintillation effects on Crab GPs and average profiles have been observed and discussed before (e.g. 1995ApJ…453..433L (54), 1990MNRAS.244…68R (74), and 2011MNRAS.410..499G (29)). Interestingly, in a recent paper, 2024ApJ…973…87D (21) discuss the possibility of using GP rate as an ISM probe, which is exactly what transpired from our observations.
Under this hypothesis and using our equation 9 and equation 4.45 from 2012hpa..book…..L (53) (see also equation 4.1 in 1986MNRAS.220…19R (75)), we obtain:
where is the timescale of the refractive scintillation, is the distance from the Earth to the screen (Fig. 10), is the standard deviation of the scattering angle, and can be interpreted as the velocity of the refractive scintillation pattern at the location of the observer (i.e. on Earth).
Additionally, the scatter broadening time is related to by the equation (e.g. 2012puas.book…..L (57)) :
where c is the speed of light.
Typically, scintillation analysis (e.g. 2014ApJ…791L..32B (8)), uses auto-correlation function (ACF) of dynamic spectrum containing scintles to calculate scintillation time and de-correlation bandwidth as respectively the full-width at 1/e and half-maximum half-width of the 2D Gaussian fitted to the correlation peak in ACF. In this analysis the fluence normalisation or number of GPs were sampled daily in 1 hour observations. Hence, long-timescale dynamic spectrum is not available, and we can only use 1D temporal dependence of the number of GPs or normalisation of the fluence distribution (). Hence, in order to be consistent with this convention, we estimate by fitting Gaussian to the peak of the normalisation of the fluence distribution (not shown here but very similar to Figure 4) as a function of time, and calculate 51 days as full width at 1/e of the peak.
In order to calculate using equation 8, the distance to the screen and are also required. The additional constraints are imposed by equation 9 (above) and the following equation (equation 3 in 2014ApJ…791L..32B (8)):
where x= defines the location of the screen, v120 km/s is the pulsar’s proper motion, v 30 km/s is the Earth’s proper motion and vs is the proper motion of the screen. We note that from now onwards we shall use the known velocities of the pulsar and Earth as the case of maximum transverse velocity, and consequently the symbol will be dropped. The combination of these three equations leads to the following 3-order equation in y:
where , , and . For the maximum scattering time 5 ms and 51 days, this equation has 2 physically acceptable solutions, , and , which lead to 0.05 pc (screen close to Earth), and 1835 pc (screen close to the pulsar).
The first solution 0.05 pc could potentially be related to the inner edge of the solar system (distance 0.01 pc), but this does not have convincing physical explanation. On the other hand, the second solution (1835 pc) places the screen close to the pulsar, and the distance matches the distance to the Perseus spiral arm of the Milky Way galaxy, where the Crab pulsar is located. Hence, the scattering screen could be placed somewhere at the edge of the Perseus arm as seen along the LoS from Earth, which was also proposed by earlier studies (e.g. 1971ApJ…166..513C (18)). Alternatively, given the accuracy of this simplistic approach, and uncertainties of the distance to the pulsar, velocities of the screen and pulsar it is possible that the scattering material corresponds to the expanding Crab Nebula surrounding the pulsar. We note that solar elongations of Crab pulsar during the observing period were between 180°and 74°- sufficiently large to make any contributions from the solar wind negligible.
[…]
| 2 | 0.030 | 1797 | 17.5 | 1851 |
| 3 | 0.037 | 1816 | 21.6 | 2279 |
| 4 | 0.040 | 1827 | 25.0 | 2640 |
| 5 | 0.050 | 1835 | 28.0 | 2957 |
The above estimates were obtained for the specific maximum scattering time 5 ms. Hence, the question arises as to what screen distance can be calculated for other scatter broadening times (i.e. different dates). We calculated for other values to between 5 and 2 ms (Tab. 2). The distance to the screen varies from about 1835 down to 1797 pc. Obviously, if increase in scatter broadening time is caused by a single plasma cloud, the distance to this cloud should remain constant within the measurements errors. However, we believe that the different distance values in Table 2 are due to measurement errors and very simplistic modelling approach. For example, orientation of all velocity vectors has been ignored, and only maximum velocity values have been used in equation 11. Simple tests show that even small changes in or corresponding to changing orientation to the Earth’s proper the Sun motion are of the similar order to the observed differences (Tab. 2) in , etc. A more complex modelling calculating for various orientations of velocity vector, screen velocities, and propagation of errors to solutions of equation 11 should be performed to further validate these statements and results. We defer more detailed modelling to future work.
6.1.3 Modelling scatter broadening time variations
We have also made an attempt to model time dependence of as the following Gaussian:
where is offset, and is the timescale of the variations. The free parameters of the fit were , and . We note that is unknown, and therefore the product was used as a free parameter of the fit.
This model can be interpreted as the LoS passing through different parts of the plasma cloud with different electron densities and amount of turbulence causing variable scattering (represented by ). The fitted curve and parameters are shown in Figure 12. Although is consistent with the earlier DM modelling, these results show that the timescale days of variations is slightly longer than fitted to DM variations (12.5 days). Fitting the data with the parameter fixed to 12.5 days (as obtained from fit to DM(t)) resulted in much worse reduced value.
Hence, the timescale of scattering the scattering is longer than the timescale of DM variations. Also, the scattering started and finished before the contribution of the plasma cloud to the total DM became measurable. It is expected that the onset of scattering becomes observable before the direct LoS is crossed by the cloud, and similar effect was observed by 2008A&A…483…13K (45).
We interpret our findings (Sections 6.1.2 and 6.1.3) using two screens model. With the variable component of the scattering caused by a screen in the expanding Crab Nebula or in the Perseus arm of the Galaxy, while the approximately constant “baseline” (2 ms) due to the second screen (not modelled here). Such a two screen model was suggested and applied by the earlier studies 1971ApJ…166..513C (18, 36, 56, 88, 31, 60). The peaks (“horns”) in scattering time at around 43 – 58 and 65 – 83 days, may be due to additional sub-structures (lumps) of turbulent plasma within the main cloud. The fit to the data excluding these structures yielded very similar results. However, we defer more complex modelling using the two screen model, and potential inclusion of the “horns” to the future work.
6.1.4 Spectral behaviour of scatter broadening
We have also made an initial attempt to measure frequency scaling of scatter broadening time using bright GPs and software package SCAMP 999https://github.com/pulsarise/SCAMP-I 2021MNRAS.504.1115O (70, 28). The band was divided into 4 sub-bands and MCMC fitting of individual GP profiles was performed in each sub-band (see example in left Figure 13 for the brightest GP from 2024-12-24), and yielded scatter broadening times () consistent with our method. Based on this, scattering spectral index (where ) was fitted to scattering time as a function of frequency (e.g. right image in Figure 13). Figure 13 shows that the package performs well on Crab GPs despite the fact that it was originally designed to work on average pulse profiles. The same fitting procedure was applied to all the datasets and yielded as a function of time. The fitting performs very well during low scattering conditions for high SNR GPs, but did not converge well when the scattering was higher (SNR lower). Several outliers and failed fits were identified and discarded, and spectral index of the scatter time as a function of date is shown in Figure 14. The mean value of the spectral index is , which is in a good agreement with the previous measurements with the MWA (e.g. 2019ApJ…874..179K (44)), and broadband measurements by 2006ARep…50..562P (72, 46). Overall the spectral index varies significantly on various dates. Especially, the SCAMP fit did not perform well during high-scattering conditions regardless if isotropic or anisotropic scattering model was used. This could be caused by either the reduced SNR or neither of the models being applicable to these high scattering conditions. Additionally, the flattening of scattering spectral index observed when scattering was high (mid-February and early March 2025) may be a result of inadequate screen model.
These results show that any measurements of scattering spectral index for Crab pulsar have to be treated with caution. Mainly because highly variable scattering conditions may impact the fitting results on a particular day (e.g. cause apparent flattening) by reducing SNR or requiring different screen models (e.g. thin vs. thick, isotropic vs. anisotropic etc.). Additionally, as noted in 2021MNRAS.504.1115O (70, 28), the spectral index measurements are strongly dependent on the fitting procedure or package (e.g. fitting in log-log space vs. original values etc.). It is clear that consistent measurements require further work on screen models and more stable fitting procedures, especially during high-scattering/low-SNR conditions. We leave these efforts for the future, and we are willing to contribute to adjustments in the SCAMP package to make it suitable for single pulse analysis.
6.2 FRB context
The possibility of extra-galactic young Crab-like pulsars being responsible for FRBs has been discussed by multiple authors (see references in Section 1). Although, the energies, distribution of times between pulses, and other arguments favour magnetars as the progenitors of majority of FRBs, in this section we revisit this possibility using our huge sample of low-frequency GPs.
6.2.1 Spectral luminosity
The maximum spectral luminosities observed in our sample are of the order of ergs/s/Hz (right plot in Fig. 8), which is consistent with earlier measurements performed typically at higher frequencies (e.g. Figure 3 in 2022NatAs…6..393N (66)). These spectral luminosities are some 3 orders of magnitude lower than the faintest pulses from repeating FRB 20200120E reported therein. By extrapolating the fitted spectral distribution of Crab GPs (right plot in Figure 8) to higher luminosities, we find that the waiting time for erg/s/Hz pulse is of the order of 66,000 years. The total number of galaxies up to redshift z8 was estimated by 2016ApJ…830…83C (14) to be . We estimate number galaxies up to redshift z=0.1 to be (the highest reported FRB redshifts are 1 2023Sci…382..294R (76)) by scaling by the ratio of co-moving volumes 101010https://www.astro.ucla.edu/~wright/CosmoCalc.html, while . We note that, scaling of by the ratio of co-moving volumes is a simplification not taking into account redshift evolution of the number of galaxies. Next, we assume number of young Crab-like pulsars per galaxy . There are no good estimates for this number and it could be as low as a 1 or as high as 100. Then a nearby FRB (z¡0.1) with spectral luminosity erg/Hz/s could be observed every few hours. Hence, FRBs similar to repeating FRB 20200120E, for which Crab-like pulsars are also considered as a potential progenitors 2021ApJ…910L..18B (6), could be relatively common. In fact, limits on number of pulses from such objects could provide estimates on the number of Crab-like pulsars.
Nevertheless, spectral luminosities of Crab GPs are very difficult to reconcile with some 15 orders higher spectral luminosities of typical one-off FRBs. Even assuming that the energy “reservoir” in Crab pulsar is sufficient to produce pulses ergs/s/Hz, the expected FRB wait times would be about several orders of magnitude longer than the Hubble time. Hence, our observations are in agreement with the current models that Crab-like pulsars may be responsible for a population of weaker repeating FRBs, but cannot produce more energetic FRBs.
6.2.2 Arrival times
Several studies 2024ApJ…973…87D (21, 54, 40) showed that the distribution of pulse separation times for the Crab pulsar follows an exponential distribution, which corresponds to a purely random, Poisson, process. Hence, the detection of a pulse does not have impact on waiting time for the next one. This differs from what is observed in repeating FRBs, in which case the distribution is better characterised by Weibull distribution resulting from some form of correlation (clustering or anti-clustering) between the subsequent pulses (e.g. in clustering, detection of a pulse increases the probability of detecting another pulse). The distribution of time separations of GPs was studied before 1995ApJ…453..433L (54, 40, 21). However, these earlier studies were performed at frequencies above 800 MHz (mainly L-band), and in some cases (e.g. 1995ApJ…453..433L (54)) only up to a time difference of about 66 s (2000 pulsar periods). Additionally, recent high-frequency (1.6 GHz) observations by 2023ApJ…959..111L (50) reported detections of multiple (up to 6) micro-bursts within a single pulse window (1 ms), largely exceeding numbers expected from a purely independent process. At our frequencies, however, micro-bursts within a pulse window (or even a few ms) are “blended” together due to large scatter broadening time. Hence, we used our giant sample of GPs to verify the distribution of pulse time separations at time scales longer than the pulsar period (verification of our merging procedure confirmed lack of pulses separated by less than 50 ms).
Figure 15 shows distributions of time between pulses for GPs with SNR10, where we only used “good data sets” (at least 3000 s of non-flagged data with at least 2000 GPs detected) to probe time differences up to 600 s calculated within each dataset only (up to 1 h observations). This ensured that the gaps between the observations or refractive scintillation did affect the analysis. The SNR10 threshold ensures that false-positives candidates, which can significantly impact the distribution, were not included. However, we verified that the SNR thresholds higher than 7 give consistent results. The resulting distribution can be very well-fitted with an exponential distribution:
resulting in (left image in Figure 15). The same fit was performed for each separate dataset, and in all cases the distribution of time between pulses was well described by an exponential distribution. The value of the fitted parameter varies between when the number of GPs detected per hour was highest (end of December 2024) down to 0.09 when the GP rate was lower (end of March 2025). The plot of fitted vs. time has similar shape to Figure 4, and has not been included here for brevity. This varying behaviour of is perfectly in line with the fact that represents number of GPs in a certain time interval (GP rate) which was changing with changing scattering conditions (as discussed in Section 5.3).
The distribution was also fitted with a Weibull distribution (right image in Figure 15) :
which reduces to the exponential distribution for shape parameter k=1, and corresponds to clustering. The fit yielded k = consistent with 1, which confirms that the distribution of pulse arrival times is well fitted with exponential distribution as expected for independent events (i.e. Poisson process).
Finally, we have also created a distribution of logarithm to the base 10 of the wait times (Fig. 16), and compared the standard deviation of the fitted log-normal distribution to those expected for a periodic process, shot noise and observed in repeating FRBs (see Table 1 in 2024OJAp….7E..43K (41)). The standard deviation of the log-normal distribution fitted to our data is between 0 expected for a purely periodic process (e.g. regular pulses from a pulsar) and 0.723 expected for shot noise. However, it is lower than observed in repeating FRBs with sufficiently large number of pulses detected, and much lower than even observed in galactic magneters (see Table 1 in 2024OJAp….7E..43K (41)). This further confirms that statistical properties of arrival times of Crab GPs are different to those observed in repeating FRBs, which also indicates differences in physical emission mechanisms and processes.
6.2.3 Limits on low-frequency FRB rate
The entire presented project was originally motivated by searches for pulses from repeating FRBs, magnetars and RRATs. Hence, during this project three repeating FRBs were observed namely: FRB20240114A (27 hours), FRB20201130A (27 hours), FRB20180301 (9 hours) and FRB20240619D (5 hours). No pulses at DM of these FRBs were detected above the detection threshold of 240 Jy ms.
Similarly, no FRB-like events other than Crab giant pulses were detected in the entire dataset (repeating FRBs, RRATs, magnetars, Crab pulsar, nearby galaxy NGC 3109, and several other fields) totalling to about 100 hours of observations with the station beam FoV0.012 srad. This negative result can be translated into an upper limit on FRB rate Rfrb in 200 – 231.25 MHz band, which is R2500 FRBs/day/sky. This limit is some 2 orders of magnitude higher than the current extrapolations of FRB rates from higher frequencies (Figure 3 in 2024PASA…41…11S (78)). Hence, the observing time and/or FoV need to increase substantially in order to routinely detect low-frequency FRBs. For example, even increasing number of beams to 10 could lead to first FRB detections after some 1000 hours of observations (about 6 weeks of non-stop observing with EDA2). However, millisecond all-sky imaging, as proposed by 2022aapr.confE…1S (83), is the best way to convert EDA2 and SKA-Low stations into “FRB machines” once the technical challenges of data capture, high-time resolution imaging, and real-time FRB searches are addressed (progress in these areas is described by 2024PASA…41…42S (81, 20)).
7 Summary and conclusions
Once more, standalone SKA-Low stations and EDA2 in particular turned out to be very flexible and productive scientific instruments for pulsar and transient sky monitoring. This work was possible thanks to the flexibility and quick turn around of the data as observations were typically processed and inspected within 24 hours from acquisition.
In this paper, we have presented a new single pulse search pipeline implemented on a single station beam from EDA2 station. The effort was driven by searches for single pulses from FRBs, RRATs, magnetars and pulsars. No pulses from known repeating or other FRBs were detected during 100 hours of observations, which resulted in upper limit for FRB rate R2500 FRBs/day/sky.
During the observing campaign nearly 100,000 sample of giant pulses (GPs) from the pulsar B0531+21 (Crab) were detected, which is to our knowledge the largest GP sample at frequencies 300 MHz. In this paper, we presented statistical analysis of this unprecedented sample of GPs. On the timescales accessible to our analysis (1 ms), the observed distribution of pulse arrival times is consistent with exponential distribution expected for purely random, Poisson process. These findings are consistent with the earlier observations that Crab GPs, unlike FRBs, do not show any signs of clustering observed in FRBs. Due to limitations of low-frequency observations, further high time resolution, high-frequency studies will be required to confirm the clustering of micro-bursts reported by 2023ApJ…959..111L (50). Our findings and the observed distribution of spectral luminosities do not exclude a possibility that extra-galactic Crab-like pulsars can produce weak repeating FRBs, but are unlikely progenitors of the entire population of FRBs. Especially, that given the observed distribution of spectral luminosities the wait times for typically bright FRBs from extra-galactic Crab-like pulsars would significantly exceed Hubble time. Hence, the FRB rates would be many orders of magnitude lower than the observed rates (depending on the fluence, from 100s to 1000s per day per sky).
The observed fluence distribution of bright (SNR10) GPs can be well characterised by a single power law with the index , and indexes and for distributions of GPs from main pulse and interpulse respectively. We observed a strong correlation (coefficient 0.9) of number of detected GPs with the measured scattering times, and when scattering is low (2 ms) a much larger number of GPs was observed than at the time of high scattering ( 5 ms). This is a result of fluence distribution normalisation being strongly modulated (correlation coefficient 0.9) by the scatter broadening time, which we recognise as the effect of refractive scintillation. The spectral index of scatter broadening was estimated to be , but the measurements were strongly affected by scattering conditions (including reduced SNR of GPs amongst other effects).
We also observed indications of correlation of DM and scatter broadening time variations (correlation coefficient 0.7), which may be overestimated due to the impact of scatter broadening on DM measurements.. Assuming that the event was caused by refractive scintillation, the estimated distance to plasma screen is of the order 1830 pc which places the screen in the Perseus arm of the Galaxy or even the Crab Nebula itself. The velocity of the scintillation pattern, km/s, is similar to the expansion velocities observed in the Crab Nebula, which further supports this possibility. Modelling of DM and variations leads to size of plasma screen pc with Gaussian distributed cloud of electron plasma of mean electron density e-cm-3, and peak electron density e-cm-3. This range of parameters is in agreement with earlier measurements, and matches electron densities in the filamentary structures of the Crab Nebula.
Our detections of Crab GPs (the faintest down to 500 Jy ms) clearly demonstrate that even a single SKA-Low station can be a powerful instrument in searching for and studying pulsars, FRBs and other fast transients. Especially, once multi-beaming capability is commissioned on the newly built SKA-Low stations (4 beams planned in Array Assembly 1 and more later). Furthermore, as discussed in 2024PASA…41…11S (78, 83, 81) the high time resolution (millisecond) imaging of the entire visible hemisphere, leveraged with the real-time FRB search pipelines (e.g. 2025arXiv250515204D (20)) can yield even one FRB detection per day. The presented analysis demonstrates the first single-beam attempt and proof of principle for such efforts on a larger scale. Once multiple beams or thousands of image pixels with the same sensitivity can be formed, the scientific output from a single station can be boosted by orders of magnitude.
Our results provide a very optimistic prognostic for SKA-Low transient science, and of what may be possible in just a few years when hundreds of SKA-Low stations become operational. They also demonstrate SKA-Low potential to study local environments of pulsars, FRBs or other fast transients using low frequencies where propagation effects are much more pronounced than at higher frequencies. Given that nearby FRBs (z¡0.1) can be similarly bright, and their DMs similar or only a few times higher than Crab pulsar’s (e.g. FRB 20200120E has DM=87.82 pc cm-3 2021ApJ…910L..18B (6)), we conclude that the huge potential of SKA-Low for wide-field and all-sky searches for FRBs and transients is only a few steps away!
Acknowledgments
EDA2 is hosted by the MWA under an agreement via the MWA External Instruments Policy. This scientific work makes use of the Murchison Radio-astronomy Observatory, operated by CSIRO. We acknowledge the Wajarri Yamatji people as the traditional owners of the Observatory site. This work was further supported by resources provided by the Pawsey Supercomputing Centre with funding from the Australian Government and the Government of Western Australia. The acquisition system was designed and purchased by INAF/Oxford University and the RX chain was design by INAF, as part of the SKA design and prototyping program. SD acknowledges the funding received from the ICRAR Summer Student program, which made parts of this work possible. We thank the anonymous referee for their helpful comments that significantly improved the quality of the manuscript.
Data availability
The data can be made available on a reasonable request in their original or final formats. The original data were complex voltages from X and Y polarisations of a station beam in 1.08 usec time resolutions.
Appendix A Modelling pulse scatter broadening
Pulse scatter broadening was modelled with a Gaussian (representing the on-set and peak of the pulse) with an exponential tail (only after the time of the pulse peak). Hence, it was mathematically described as:
where is the Gaussian profile used to describe the pulse peak:
and t is the time, is standard deviation of the Gaussian profile, is the time of the peak flux density , and is the scatter broadening time. An example of a Crab giant pulse fitted with the above equations is shown in the right panel of Figure 2.
We also tested more standard description where the Gaussian was convolved with an exponential function for times t above :
However, both parametrisations yielded very similar results.
Appendix B DM verification by maximising peak flux of GPs
It was noticed that high scattering may impact accuracy of timing-based DM measurements (mainly by shifting and changing shape of the average pulse profiles). Therefore, in order to verify the DM measurements obtained from timing we also measured DM using very bright (SNR20) individual GPs. We tested the following methods of finding the optimal DM by : (a) maximising the slope of the leading edge between and of the peak, (b) maximising the slope of the entire leading edge, (c) maximising the peak flux density (i.e. SNR) and (d) maximising the rise time (difference between the peak time and start of the leading edge). The first two methods (a,b) were significantly affected by the scatter-broadening which led to large errors and variations of the resulting DM during high-scattering period. Therefore, we rely on method (c), which is also the typically used by single pulse search software packages finding DM maximising SNR/peak flux. The method (d) gives very similar results to method (c), but was also less reliable during high-scattering times. The comparison of DM measured with standard pulsar timing and by maximising peak flux density (method (c)) is shown in Figure 17. The errors in method (c) are much larger than from timing analysis during high scattering (between 10 and 28 February 2025) mainly because method (c) relies on single GPs (highly scattered and low-SNR at times), while timing analysis uses entire (1-hour) datasets. However, despite signs of systematic offset (especially between our measurements and Jodrell Bank data) the two methods are consistent, and indicate that the increase in DM, and its positive correlation with scattering time are real effects (i.e. not artefacts of analysis caused by the scatter broadening). This comparison, as well as comparison with DM measured by Jodrell Bank gave us confidence that our DM measurements are to within a factor of 2 accurate.